# Classification of Semigraphical Translators

### v1

@@ Abstract.
F. Martín was partially supported by the grant PID2020-116126-I00 funded by MICIU/AEI/ 10.13039/501100011033 and by the IMAG-Maria de Maeztu grant CEX2020-001105-M funded by MICIU/AEI/10.13039/501100011033.
@@ 1. Introduction
A surface $\Sigma$ in $\mathbf{R}^{3}$ is called a translator (for mean curvature flow) with velocity $\mathbf{v}$ if
$\Sigma\mapsto\Sigma_{t}:=\Sigma+t\mathbf{v}$
is a mean curvature flow, i.e., if the normal velocity at each point is equal to the mean curvature vector: $\vec{H}=\mathbf{v}^{\perp}$. By rotating and scaling, it suffices (for nonzero velocities) to consider the case $\mathbf{v}=-\mathbf{e}_{3}$. In that case, we refer to $\Sigma$ simply as a translator. Ilmanen [ilmanen-1994] observed that $\Sigma$ is a translator if and only if it is minimal with respect to the translator metric $g_{ij}(x,y,z)=e^{-z}\delta_{ij}$.
If the surface $\Sigma$ is the graph of a function $u$ over a region of the $(x,y)$-plane, then the translator equation $\vec{H}=(-\mathbf{e}_{3})^{\perp}$ is equivalent to
@@
A smooth, connected, complete properly embedded translator $\Sigma$ is called semigraphical if it contains a non-empty, discrete collection $\{L_{n}\}_{n}$ of vertical lines such that $M\setminus\bigcup_{n}L_{n}$ is a graph $z=z(x,y)$. This notion was introduced by Hoffman-Martín-White in [tridents], which included a partial classification. Building on their work, we complete the classification of semigraphical translators and prove
@@
A semigraphical translator $\Sigma$ in $\mathbf{R}^{3}$ is one of the following:
@@
Theorem 1, minus Assertions (i), (ii), and (iii), is the same as Theorem 34 of [tridents], except that the list in Theorem 34 included a sixth item, sometimes called a “yeti”. In [tridents], yetis were conjectured not to exist. Asssertions (i), (ii), and (iii) assert that, for the allowed values of the parameters, the indicated surfaces exist and are unique. Existence and uniqueness for the tridents was proved in [tridents], and existence and uniqueness for the other surfaces was proved in [scherk-like], except that uniqueness of pitchforks and uniqueness of translating helicoids were left as conjectures. In this paper, we complete the proof of Theorem 1 by proving those three conjectures: uniqueness for pitchforks (§3) and translating helicoids (§4), and nonexistence of the yeti (§5).
@@
[…]
@@
Pitchforks are simply connected and have entropy $3$. They might therefore arise as blowups of mean curvature flows of surfaces that are initially closed and smoothly immersed. (By recent work [bamler-kleiner] of Bamler and Kleiner, they cannot arise as blowups if the initial surface is embedded.) All semigraphical translators other than pitchforks and tridents have cubic area growth, and therefore infinite entropy, and hence cannot arise as such blowups. Tridents have finite entropy but infinite genus, and thus they cannot arise as blowups at the first singular time.
@@
Figure 1. Left: A fundamental piece of the pitchfork of width $\pi$. Right: The whole surface, obtained from the fundamental piece by a $180^{\rm o}$ rotation around the $z$-axis.)
@@
There are interesting translators that are neither graphical nor semigraphical. Examples include: helicoid-like translators that contain a vertical line and are invariant under a screw motion about that line [helicoidal], surfaces obtained by gluing [davila, nguyen-reapers, nguyen-doubly], rotationally-invariant annuli [clutterbuck], and a large family [annuli] of annuli contained in slabs (and therefore not rotationally invariant).
The proofs in this paper use Morse-Radó theory [morse-rado], which studies critical points and level sets for a class of functions arising in minimal surface theory. (Recall that translators are minimal surfaces with respect to the translator metric.) The uniqueness proofs use an angular maximum principle that may be of independent interest. The proof of non-existence of the yeti involves comparison with an appropriately placed grim reaper surface, using a technique first employed by Chini [chini].
@@ 2. An Angular Maximum Principle
In this section, we prove a version of the maximum principle that will be used to prove the uniqueness of pitchforks and helicoidal translators.
@@
$(u_{2}-u_{1})|\partial\Omega=d$
for some constant $d$. Suppose also that
$\inf_{(x,y)\in\Omega\,x\geq x_{0}}\nu_{i}(x,y)\cdot\mathbf{e}_{2}>\epsilon>0$
@@
By translating, we can assume that $b=0$.
Suppose, contrary to the theorem, that $u_{2}-u_{1}$ is not constant. By relabelling, we may assume that there are points in $\Omega$ where $u_{2}-u_{1}>d$.
@@
$\mu:=\max_{\Omega\cap\{x\leq x_{0}\}}(u_{2}-u_{1})<\sup(u_{2}-u_{1}).$
By replacing $\Omega$ by $\{(x,y)\in\Omega:u_{2}-u_{1}\geq\mu\}$ and by replacing $u_{2}$ by $u_{2}-\mu-d$, we can assume that $\Omega$ lies in
$[x_{0},\infty)\times(0,c),$
that $u_{2}=u_{1}$ on $\partial\Omega$, and that $u_{2}\geq u_{1}$ on $\Omega$. By the strong maximum principle, $u_{2}>u_{1}$ on $\Omega$.
By translating, we can assume that $x_{0}$ is as large as we like. In particular, we can assume that
@@
Since $x_{0}>0$, we have well-defined polar coordinates (with $20\leq\theta<\pi/$ 2) on $\overline{\Omega}$.
@@
Now we define a function $\omega:\Omega\to\mathbf{R}$ as follows. If $p\in\Omega$, let $J(p)$ be the connected component of
@@
Since $u_{1}=u_{2}$ at the endpoints of $J(p)$ and since $\partial u_{i}/\partial\theta<0$ at all points of $J(p)$, we see that $u_{1}$ and $u_{2}$ map $J(p)$ diffeomorphically to the same interval. Thus there is a unique $q\in J(p)$ such that $u_{2}(q)=u_{1}(p)$. We define
$\omega(p)=\theta(q)-\theta(p).$
@@
Note that if $x(p)>0$ and $0\leq y(p)\leq c$, then
$\theta(p)\geq 0$
and
$r(p)\sin\theta(p)\leq c,$
so $\theta(p)\leq\arcsin(r(p)^{-1}c)$.
It follows that for every $p\in\Omega$,
$0<\omega(p)<\arcsin(r(p)^{-1}c).$
@@ 3. Uniqueness for Pitchforks
$f(x,w)$ ==$\infty$,= - ∞ ,
$f(x,0)$ =$=\begin{cases}-\infty&\text{if $x<0$},\\ \infty&\text{if $x>0$},\end{cases}$
@@
Here, $M$ is half of a pitchfork, as in Fig. 1 (viewed from a point in the first octant). The other half, also shown in Fig. 1, is obtained by rotating $M$ by $\pi$ about $Z$. The hypothesis $w\geq\pi$ is necessary: if $w<\pi$, then no such $f$ exists [scherk-like]*Theorem 12.1.
@@
Existence was proved in [scherk-like]*§12, so it suffices to prove uniqueness. To prove uniqueness, it will be convenient to add two points $+\widehat{\infty}$ and $-\widehat{\infty}$ at infinity to $\mathbf{R}^{2}$ such that, for $p_{i}\in\mathbf{R}^{2}$,
@@
(*) $f_{1}(p)-f_{2}(p+v)$
has a critical point at $q$. Note that a critical point of (*3) is equivalent to a point of tangency between the graph of $f_{2}(p+v)$ and the foliation whose leaves are vertical translates of $\operatorname{graph}(f_{1})$. It follows, by [morse-rado]*Corollary 40, that the function (*3) has a critical point for all $v$ sufficiently close to $q^{\prime}-q$. (Indeed, the number of critical points, counting multiplicity, of (*3) is a lower-semicontinuous function of $v$.) In particular, there exists such a $v$ with $y(v)\neq 0$. ∎
@@
where $b=w-y(v)$. Note that $0<b<w$, and that $g$ has boundary values
@@
Let $C$ be a component of the zero set of $g$ that contains the critical point $p$. Then $C$ has the structure of a network in which $p$ is a node, and in which the valence of each node is an even number $\geq 4$. (See, for example, [colding-minicozzi]*Theorem 7.3.) By the maximum principle, $C$ contains no closed loops, so it is a tree with at least $4$ ends. From the boundary values of $g$, we see that each end tends to $-\widehat{\infty}$, to $+\widehat{\infty}$, or to $(0,0)$.
@@
We claim that there is a $\tilde{x}<0$ such that
@@
00 0 $\geq\frac{\partial g}{\partial y}(x_{i},y_{i})$
@@
since $\psi$ is strictly decreasing, a contradiction. Thus, there is an $\tilde{x}$ for which (4) holds. Hence, for $x\leq\tilde{x}$, there is at most one $y$ such that $(x,y)\in C$. Thus, at most one end of $C$ tends to $-\widehat{\infty}$.
Since at most one end of $C$ tends to $-\widehat{\infty}$ and at most one end tends to $(0,0)$, at least two ends must tend to $+\widehat{\infty}$. Hence, $C$ contains two embedded curves $\alpha$ and $\beta$ that join $p$ to $+\widehat{\infty}$ and that are disjoint except at $p$. Let $\Omega$ be the open region bounded by $\alpha\cup\beta$. Then $g=0$ on $\partial\Omega$ and $g$ does not vanish anywhere on $\Omega$. But, by (3) and the angular maximum principle (Theorem 2), $g\equiv 0$ on $\Omega$, a contradiction. ∎
@@ 4. Uniqueness for Translating Helicoids
$f(x,0)$ =$=\begin{cases}\infty&\text{if $x<0$},\\ -\infty&\text{if $x>0$},\end{cases}$
$f(x,w)$ =$=\begin{cases}-\infty&\text{if $x<a$},\\ \infty&\text{if $x>a$},\end{cases}$
@@
    $\nu(p)$ converges to $\mathbf{e}_{2}$ as $p\to-\widehat{\infty}$ and to $-\mathbf{e}_{2}$ as $p\to+\widehat{\infty}$.
@@
Let $C$ be the connected component of $g^{-1}(0)$ containing $p$. Exactly as in the proof of Theorem 3, $C$ is a tree with at least four ends, and each end tends to $(0,0)$, $(a,b)$, $-\widehat{\infty}$, or $+\widehat{\infty}$. Also, as in that proof, at most one end tends to each of those points. (In the cases of $\pm\widehat{\infty}$, this is by Property (P2) and the angular maximum principle, Theorem 2.) Thus, exactly one end tends to each of those points. Consequently, $C$ contains a curve $C^{\prime}$ that joins $(0,0)$ to $(a,b)$.
@@
Since $C$ has an end tending to $+\widehat{\infty}$, there are points in $U$ at which $g=0$. By the strong maximum principle, there are also points at which $g>0$. Thus the set

### v2

@@ Abstract.
F. Martín was partially supported by the grants PID2020-116126-I00 and PID2024-156031NB-I00 funded by MICIU/AEI/ 10.13039/501100011033 and by the IMAG-Maria de Maeztu grant CEX2020-001105-M funded by MICIU/AEI/10.13039/501100011033.
@@ 1. Introduction
A surface $M$ in $\mathbf{R}^{3}$ is called a translator (for mean curvature flow) with velocity $\mathbf{v}$ if
$M\mapsto M_{t}:=M+t\mathbf{v}$
is a mean curvature flow, i.e., if the normal velocity at each point is equal to the mean curvature vector: $\vec{H}=\mathbf{v}^{\perp}$. By rotating and scaling, it suffices (for nonzero velocities) to consider the case $\mathbf{v}=-\mathbf{e}_{3}$. In that case, we refer to $M$ simply as a translator. Ilmanen [ilmanen-1994] observed that $M$ is a translator if and only if it is minimal with respect to the translator metric $g_{ij}(x,y,z)=e^{-z}\delta_{ij}$.
If the surface $M$ is the graph of a function $u$ over a region of the $(x,y)$-plane, then the translator equation $\vec{H}=(-\mathbf{e}_{3})^{\perp}$ is equivalent to
@@
A smooth, connected, complete properly embedded translator $M$ is called semigraphical if it contains a non-empty, discrete collection $\{L_{n}\}_{n}$ of vertical lines such that $M\setminus\bigcup_{n}L_{n}$ is a graph $z=z(x,y)$. This notion was introduced by Hoffman-Martín-White in [tridents], which included a partial classification. Building on their work, we complete the classification of semigraphical translators and prove
@@
A semigraphical translator $M$ in $\mathbf{R}^{3}$ is one of the following:
@@
Theorem 1, minus Assertions (i), (ii), and (iii), is the same as Theorem 34 of [tridents], except that the list in Theorem 34 included a sixth item, sometimes referred to as a “yeti” (or “abominable flowman”). In [tridents], yetis were conjectured not to exist. Assertions (i), (ii), and (iii) assert that, for the allowed values of the parameters, the indicated surfaces exist and are unique. Existence and uniqueness for the tridents was proved in [tridents], and existence and uniqueness for the other surfaces was proved in [scherk-like], except that uniqueness of pitchforks and uniqueness of translating helicoids were left as conjectures.
In this paper, we complete the proof of Theorem 1 by proving those three conjectures: uniqueness for pitchforks (§3) and translating helicoids (§4), and non-existence of the yeti (§5).
@@
    for Scherk translators, a parallelogram $\mathcal{P}(\alpha,w,L)$ of base angle $\alpha\in(0,\pi)$, width $w$, and unique length $L(\alpha,w)$, with boundary values $\pm\infty$ on the pairs of opposite sides;
@@
Pitchforks are simply connected and have entropy $3$. They might therefore arise as blowups of mean curvature flows of surfaces that are initially closed and smoothly immersed. (By recent work [bamler-kleiner] of Bamler and Kleiner, they cannot arise as blowups if the initial surface is embedded.) All semigraphical translators other than pitchforks and tridents have cubic area growth, and therefore infinite entropy, hence they cannot arise as such blowups. Tridents have finite entropy but infinite genus, and thus they cannot arise as blowups at the first singular time.
@@
Figure 1. Left: A fundamental piece of the pitchfork of width $\pi$. Right: The whole surface, obtained from the fundamental piece by a $180^{\circ}$ rotation around the $z$-axis.
@@
There are interesting translators that are neither graphical nor semigraphical. Examples include: helicoid-like translators that contain a vertical line and are invariant under a screw motion about that line [helicoidal], surfaces obtained by gluing [davila, nguyen-reapers, nguyen-doubly, smith], rotationally invariant annuli [clutterbuck], and a large family [annuli] of annuli contained in slabs (and therefore not rotationally invariant).
Let us indicate the main ideas of the paper by examining the geometry of a potential yeti, in contrast with that of the pitchfork and the helicoid of Figures 1 and 2. It is not hard to see that a “yeti” translator $\mathcal{Y}$, if it existed, would necessarily have very exotic properties. Notably, the Gauss map of $\mathcal{Y}$ restricted to any half-strip would come arbitrarily close to the equator of the upper hemisphere, and the gradient of the graphing function $u$ would be unbounded along every $x$-ray. Moreover, the entropy of $\mathcal{Y}$ would be either infinite or (by standard considerations, cf. [gama-martin-moller-2025]) an integer; it is conjectured that the latter is impossible for non-collapsed translators. These conditions would force very complicated geometric behavior, which we rule out using maximum principle techniques.
The arguments involved in proving Theorem 1 use Morse-Radó theory [morse-rado], which studies critical points and level sets for a class of functions arising in minimal surface theory. (Recall that translators are minimal surfaces with respect to the translator metric.) The uniqueness proofs use an angular maximum principle that may be of independent interest. The proof of non-existence of the yeti involves comparison with a family of appropriately placed grim reaper surfaces, inspired by a technique first employed by Chini [chini]. This sliding-maximum principle allows us to derive a gradient bound for the solution $u$ in a half-space. Then, an asymptotic limit of $\mathcal{Y}$ would be a complete graphical translator over the half-space, violating the classification Theorem of [graphs-himw].
@@ 2. An Angular Maximum Principle
In this section, we prove a version of the maximum principle that will be used to establish the uniqueness of pitchforks and helicoidal translators.
@@
$(u_{2}-u_{1})|_{\partial\Omega}=d$
for some constant $d$. Suppose also that their normal vectors $\nu_{i}:=\frac{(-Du_{i},1)}{\sqrt{1+|Du_{i}|^{2}}}$ satisfy
$\inf_{(x,y)\in\Omega\cap\{x\geq x_{0}\}}\nu_{i}(x,y)\cdot\mathbf{e}_{2}>\epsilon>0$
@@
After translating the domain and replacing $u_{2}$ by $u_{2}-d$, we can assume that $b=0$ and $d=0$.
Suppose, contrary to the theorem, that $u_{2}-u_{1}$ is not constant. By relabelling, we may assume that there are points in $\Omega$ where $u_{2}>u_{1}$.
@@
$\mu:=\sup_{\overline{\Omega\cap\{x\leq x_{0}\}}}(u_{2}-u_{1})<\sup(u_{2}-u_{1}).$
By replacing $\Omega$ by $\{(x,y)\in\Omega:u_{2}-u_{1}>\mu\}$ and then replacing $u_{2}$ by $u_{2}-\mu$, we can assume that $\Omega$ lies in
$x_{0},\infty)\times(0,c),$
that $u_{2}=u_{1}$ on $\partial\Omega$, and that $u_{2}\geq u_{1}$ on $\Omega$. By the strong maximum principle, $u_{2}>u_{1}$ in $\Omega$.
By translating the above configuration in the $x$-direction, we can assume that $x_{0}$ is as large as we like. In particular, we can assume that
@@
Since $x_{0}>0$, we have well-defined polar coordinates $(r,\theta)$ on $\overline{\Omega}$, with $0\leq\theta<\pi/2$. Moreover, $0\leq r(p)\sin\theta(p)=y(p)\leq c$ for any $p\in\overline{\Omega}$.
@@
Now we define a function $\omega:\Omega\to\mathbf{R}$ as follows. If $p\in\Omega$, let the circular arc $\mathcal{C}_{p}$ be the connected component of
@@
Since $u_{1}=u_{2}$ at the endpoints of $\mathcal{C}_{p}$ and since $\partial u_{i}/\partial\theta<0$ at all points of $\mathcal{C}_{p}$, we see that $u_{1}$ and $u_{2}$ map $\mathcal{C}_{p}$ diffeomorphically to the same interval. Thus there is a unique $q\in\mathcal{C}_{p}$ such that $u_{2}(q)=u_{1}(p)$. We define
(2) $\omega(p)=\theta(q)-\theta(p).$
@@
For $p,q$ as in (2), we have $x(p),x(q)>0$ and $0\leq y(p),y(q)\leq c$, thus
$\theta(p)\geq 0,\qquad r(p)=r(q).$
We conclude that for every $p\in\Omega$ with $\omega(p)=\theta(q)-\theta(p)$,
$0<\omega(p)$ $\leq\theta(q)=\arcsin(r(q)^{-1}c)=\arcsin(r(p)^{-1}c).$
@@ 3. Uniqueness for Pitchforks
$f(x,w)$ $=-\infty,$
$f(x,0)$ $=\begin{cases}-\infty&\text{if $x<0$},\\ \infty&\text{if $x>0$},\end{cases}$
@@
Here, $M$ is half of a pitchfork, as in Fig. 1 (viewed from a point in the first octant). The other half, also shown in Fig. 1, is obtained by rotating $M$ by $\pi$ about $Z$.
Remark 4.
The hypothesis that $M$ is a smooth manifold-with-boundary is not necessary; it is implied by the other hypotheses. See Proposition 11.
The hypothesis $w\geq\pi$ is necessary: if $w<\pi$, then no such $f$ exists [scherk-like]*Theorem 12.1.
@@
Existence was proved in [scherk-like]*§12, so it suffices to prove uniqueness. To prove uniqueness, it will be convenient to add two points $\widehat{\infty}$ and $\widehat{\infty}$ at infinity to $\mathbf{R}^{2}$ such that, for $p_{i}\in\mathbf{R}^{2}$,
@@
($\star$) $f_{1}(p)-f_{2}(p+v)$
has a critical point at $q$. Note that a critical point of ($\star$3) is equivalent to a point of tangency between the graph of $f_{2}(p+v)$ and the foliation whose leaves are vertical translates of $\operatorname{graph}(f_{1})$. It follows, by [morse-rado]*Corollary 40, that the function ($\star$3) has a critical point for all $v$ sufficiently close to $q^{\prime}-q$. (Indeed, the number of critical points, counting multiplicity, of ($\star$3) is a lower-semicontinuous function of $v$.) In particular, there exists such a $v$ with $y(v)\neq 0$. ∎
@@
where b=wy(v)b=w-y(v). Note that $0<b<w$, and that $g$ has boundary values
@@
Let $C$ be a component of the zero set of $g$ that contains the critical point $p$. Then $C$ has the structure of a network in which $p$ is a node, and in which the valence of each node is an even number $\geq 4$. (See, for example, [colding-minicozzi]*Theorem 7.3.) By the maximum principle, $C$ contains no closed loops, so it is a tree with at least $4$ ends. From the boundary values of $g$, we see that each end tends to $\widehat{\infty}$, to $\widehat{\infty}$, or to $(0,0)$.
@@
We claim that there is an $\tilde{x}<0$ such that
@@
$0$ $\geq\frac{\partial g}{\partial y}(x_{i},y_{i})$
@@
since $\psi$ is strictly decreasing, a contradiction. Thus, there is an $\tilde{x}$ for which (5) holds. Hence, for $x\leq\tilde{x}$, there is at most one $y$ such that $(x,y)\in C$. Thus, at most one end of $C$ tends to $\widehat{\infty}$.
Since at most one end of $C$ tends to $\widehat{\infty}$ and at most one end tends to $(0,0)$, at least two ends must tend to $\widehat{\infty}$. Hence, $C$ contains two embedded curves $\alpha$ and $\beta$ that join $p$ to $\widehat{\infty}$ and that are disjoint except at $p$. Let $\Omega$ be the open region bounded by $\alpha\cup\beta$. Then $g=0$ on $\partial\Omega$ and $g$ does not vanish anywhere on $\Omega$. But, by (4) and the angular maximum principle (Theorem 2), $g\equiv 0$ on $\Omega$, a contradiction. ∎
@@ 4. Uniqueness for Translating Helicoids
$f(x,0)$ $=\begin{cases}\infty&\text{if $x<0$},\\ -\infty&\text{if $x>0$},\end{cases}$
$f(x,w)$ $=\begin{cases}-\infty&\text{if $x<a$},\\ \infty&\text{if $x>a$},\end{cases}$
@@
Remark 6.
The hypothesis that $M$ is a smooth manifold-with-boundary is not necessary; it is implied by the other hypotheses. See Proposition 11.
@@
    $\nu(p)$ converges to $\mathbf{e}_{2}$ as $p\to-\widehat{\infty}$ and to $\mathbf{e}_{2}$ as $p\to+\widehat{\infty}$.
@@
Let $C$ be the connected component of $g^{-1}(0)$ containing $p$. Exactly as in the proof of Theorem 3, $C$ is a tree with at least four ends, and each end tends to $(0,0)$, $(a,b)$, $\widehat{\infty}$, or $\widehat{\infty}$. Also, as in that proof, at most one end tends to each of those points. (In the cases of $\pm\widehat{\infty}$, this is by Property (P2) and the angular maximum principle, Theorem 2.) Thus, exactly one end tends to each of those points. Consequently, $C$ contains a curve $C^{\prime}$ that joins $(0,0)$ to $(a,b)$.
@@
Since $C$ has an end tending to $\widehat{\infty}$, there are points in $U$ at which $g=0$. By the strong maximum principle, there are also points at which $g>0$. Thus the set